OpenStudy (anonymous):

What is the equation of the line which includes points (1, 1) and (1, 5)?

OpenStudy (anonymous):

First you need to find the slope, so subtract the y values and x values such that:\[(y_2-y_1)/(x_2-x_1)=m\] so that m is the slope. Then use the point slope formula \[y-y_1=m(x-x_1)\] where y_1 and x_1 are the coordinates of any point on the line. and solve for y.

OpenStudy (anonymous):

\[m = \frac{y_{1} - y_{2}}{x_{1}- x_{2}} = \frac{5 - 1}{1 - 1} = \frac40\] Therefore, this is a vertical line and therefore undefined. A vertical line is written in the form x = ? x = 1 since it goes through x.

OpenStudy (anonymous):

x= 1 because only a vertical line goes through both of those 2 points